Effect of Financing on Earnings Per Share

Three different plans for financing an $4,200,000 corporation are under consideration by its organizers. Under each of the following plans, the securities will be issued at their par or face amount, and the income tax rate is estimated at 40% of income:

Plan 1 Plan 2 Plan 3
10% Bonds _ _ $2,100,000
Preferred 10% stock, $40 par _ $2,100,000 1,050,000
Common stock, $4.2 par $4,200,000 2,100,000 1,050,000
Total $ 4,200,000 $ 4,200,000 $ 4,200,000

Required:

1. Determine the earnings per share of common stock for each plan, assuming that the income before bond interest and income tax is $8,400,000. Enter answers in dollars and cents, rounding to two decimal places.

Earnings Per Share on Common Stock
Plan 1 $
Plan 2
Plan 3

2. Determine the earnings per share of common stock for each plan, assuming that the income before bond interest and income tax is $3,990,000. Enter answers in dollars and cents, rounding to two decimal places.

Earnings Per Share on Common Stock
Plan 1 $
Plan 2
Plan 3

3. The principal   of Plan 1 is that it involves only the issuance of common stock, which does not require a periodic interest payment or return of principal, and a payment of preferred dividends   required.
Answer #1

Requirement 1 EBIT Interest EBT Tax EAT Preferred Dividend EPS Equity earnings G=E-F $ 5,040,000 B F I=G/H A 8,400,000 $ CEA-